OpenStudy (anonymous):

a car travels at an average speed of 48 miles per hour. how many miles does it travel in 5 hrs and 45 min'

OpenStudy (anonymous):

\[\frac{ 48 miles }{ 1 hour } \times 5\frac{ 3 }{ 4 } hours\] = 48 * 23/4

OpenStudy (anonymous):

48 times 23 divided by 4 equals 276